OpenStudy (diamondanon2):

Perform the required operations on the following functions. Given: f(x) = 3 - x; g(x) = -2x Find g[f(x)].

OpenStudy (diamondanon2):

The answer choices are 2x-6 -x-6 2x+3

OpenStudy (diamondanon2):

To me it seems we have to input values, but I do not know how. If I had to guess I would say -2x[3-x{x)]

jhonyy9 (jhonyy9):

given f(x) = 3-x and g(x) = -2x so to find g(f(x)) just substitute the f(x) in place of x inside the equation of g(x) ok. ?

OpenStudy (diamondanon2):

liike I did?

OpenStudy (diamondanon2):

or did I do it wrong

jhonyy9 (jhonyy9):

not is right because you substitute in place of x so this x will be eliminate and you get g(f(x)) = -2(f(x)) so how you wrote there above are to much x-es there ok. ?

OpenStudy (diamondanon2):

okay so what is the correct version?

jhonyy9 (jhonyy9):

g(f(x)) = -2(3-x) = ?

jhonyy9 (jhonyy9):

what you dont know ?

OpenStudy (diamondanon2):

so it is 2x+3

jhonyy9 (jhonyy9):

why +3 ?

OpenStudy (diamondanon2):

the two negative would cancel out right?

jhonyy9 (jhonyy9):

-2*3 = ?

OpenStudy (diamondanon2):

-6

OpenStudy (diamondanon2):

oh okay so 2x-6

jhonyy9 (jhonyy9):

exactly

OpenStudy (diamondanon2):

Thanks a bunch!

jhonyy9 (jhonyy9):

was my pleasure but do you understand it now sure ?

OpenStudy (diamondanon2):

Yes i believe so. But I do have one more.

jhonyy9 (jhonyy9):

ok just post it on the open questions column and tag me there please